OpenStudy (anonymous):

SOMEONE PPLLZZZ HELP!!!!!!! what it the value of the expression √153-√17+√68

OpenStudy (anonymous):

12.3693168769-4.12310562562+8.24621125124

OpenStudy (anonymous):

Simplify: 12.37-4.12+8.25

OpenStudy (anonymous):

12.37-4.12= 8.25+8.25=16.50.

OpenStudy (anonymous):

A. √17 b. 4√17 C. 2√51 D. 8√51 those are the options and thats not how you do it i think

OpenStudy (anonymous):

We have : \[\sqrt{153}-\sqrt{17}+\sqrt{68}=\sqrt{9\times17}-\sqrt{17}+\sqrt{4\times17} \\=3\sqrt{17}-\sqrt{17}+2\sqrt{17} \\=4\sqrt{17} \]

OpenStudy (anonymous):

thnx noura can you help me with accouple more

OpenStudy (anonymous):

Of course !

OpenStudy (anonymous):

ok Enrique has two rectangular dive flags. The flags are similar and have dimensions, in inches, as shown in the figure below. the large one has a width of 3x and the length of 4x-8 the small one has a with of 15 and the length of 16 what is the value in inches of x?

OpenStudy (anonymous):

A. 6 B. 9 C. 10 D.12

OpenStudy (anonymous):

i think its 9 but im not 100%